please help me solve this equation:4x^2+2x=0

Answer by stanbon(75887)

please help me solve this equation:
4x^2+2x=0
---
Factor:
2x(2x+1) = 0
x = 0 or x = -1/2
-----------------------------
Cheers,
Stan H.
=================